OpenStudy (anonymous):

what is the solution of the system? Use substitution. y = x + 6 y = 2x HELP PLEASE!!

hartnn (hartnn):

substitute y=2x in 1st equation 2x = x+6 can u solve this for x ?

OpenStudy (binary3i):

y= x+6 y=2x so x+6=2x x=6 put it into one of you equation y=2x y=2*6 y=12

hartnn (hartnn):

do not give entire solution, its against code of conduct, instead make the asker solve it.

hartnn (hartnn):

@binary3i

OpenStudy (binary3i):

sorry boss

OpenStudy (anonymous):

it better for me to see the hole thing then step by step @hartnn thank you@binary3i

hartnn (hartnn):

trust me, you won't learn like that.

OpenStudy (anonymous):

beleave it or not its help me so far i lose intrest fast

hartnn (hartnn):

you'll learn only if u do it by yourself. but the final choice is yours...

OpenStudy (binary3i):

what @hartnn says is correct. you'r going to ask again same kind of question.lol

OpenStudy (anonymous):

lol never know

hartnn (hartnn):

^ exactly learn it once and for all....

OpenStudy (binary3i):

yo

OpenStudy (anonymous):

okay thx guys